OpenStudy (anonymous):

From a balloon 1191 feet high, the angle of depression to the ranger headquarters is 72°54'. How far is the headquarters from a point on the ground directly below the balloon (to the nearest foot)?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[1191\text{ Cot}\left[\left(72+\frac{54}{60}\right){}^{\circ}\right]=366.399 \text { ft} \]

OpenStudy (anonymous):

ohhhh okay(:

OpenStudy (anonymous):

Angle of depression is the same as the angle of elevation from the point of view of the ranger headquarters.